# DefVINS: Visual–Inertial Odometry for Deformable Scenes

### v1

@@ IV-C Visual reprojection term
$\mathcal{L}_{\mathrm{vision}}^{t}=\sum_{k}\left\|\mathbf{z}_{k}-\pi\!\left(\mathtt{T}_{t},\mathbf{X}_{k}\right)\right\|_{\boldsymbol{\Sigma}_{k}},$ (10)
where $\mathbf{z}_{k}$ denotes the image measurement of feature $k$, $\mathbf{X}_{k}$ its corresponding 3D point, $\pi(\cdot)$ the camera projection model, and $\boldsymbol{\Sigma}_{k}$ the associated measurement covariance [28].
IV-D Non-rigid terms
@@
Real environments may exhibit mild non-rigid motion, such as cloth or cable deformations. To model these effects without resorting to dense representations, a lightweight deformation graph built from long-term feature tracks is adopted. Each track $i$ defines a graph node with 3D position $\mathbf{x}_{i}^{t}$ at keyframe $t$. Two nodes are connected by an edge $(i,j)$ if their distance in the reference keyframe is below a spatial threshold. This provides a simple neighborhood structure that describes how nearby points of the scene relate to each other. The reference distance between nodes is defined as
$d_{ij}^{0}=\bigl\lVert\mathbf{x}_{i}^{0}-\mathbf{x}_{j}^{0}\bigr\rVert.$ (11)
The full non–rigid regularization is built up by means of specific terms which will be explained below.
@@
$\mathcal{O}=\left[\begin{array}[]{cccccccccc}-\mathbf{J}_{r}^{-1}(\boldsymbol{\phi})\mathtt{R}_{j}^{\top}\mathtt{R}_{i}&\mathbf{0}&\mathbf{0}&\mathbf{J}_{r}^{-1}(\boldsymbol{\phi})&\mathbf{0}&\mathbf{0}&\mathbf{0}&\mathbf{0}&\mathbf{0}&\mathbf{0}\\[3.01389pt] \left(\mathtt{R}_{i}^{\top}(\mathbf{v}_{j}-\mathbf{v}_{i}-\|\mathbf{g}\|\hat{\mathbf{g}}\Delta t_{ij})\right)^{\wedge}&-\mathtt{R}_{i}^{\top}&\mathbf{0}&\mathbf{0}&\mathtt{R}_{i}^{\top}&\mathbf{0}&\mathbf{0}&-\mathtt{R}_{i}^{\top}\Delta t_{ij}&-\|\mathbf{g}\|\mathtt{R}_{i}^{\top}\Delta t_{ij}&\mathbf{0}\\[3.01389pt] \left(\mathtt{R}_{i}^{\top}(\mathbf{p}_{j}-\mathbf{p}_{i}-\mathbf{v}_{i}\Delta t_{ij}-\tfrac{1}{2}\|\mathbf{g}\|\hat{\mathbf{g}}\Delta t_{ij}^{2})\right)^{\wedge}&-\mathtt{R}_{i}^{\top}\Delta t_{ij}&-\mathtt{R}_{i}^{\top}&\mathbf{0}&\mathbf{0}&\mathtt{R}_{i}^{\top}&\mathbf{0}&-\mathtt{R}_{i}^{\top}\tfrac{1}{2}\Delta t_{ij}^{2}&-\tfrac{1}{2}\|\mathbf{g}\|\mathtt{R}_{i}^{\top}\Delta t_{ij}^{2}&\mathbf{0}\\[3.01389pt] \mathbf{H}_{\mathtt{R}_{i}}^{\text{vis}}&\mathbf{0}&\mathbf{H}_{\mathbf{p}_{i}}^{\text{vis}}&\mathbf{H}_{\mathtt{R}_{j}}^{\text{vis}}&\mathbf{0}&\mathbf{H}_{\mathbf{p}_{j}}^{\text{vis}}&\mathbf{0}&\mathbf{0}&\mathbf{0}&\mathbf{H}_{\mathrm{NR}}^{\text{vis}}\\[3.01389pt] \mathbf{0}&\mathbf{0}&\mathbf{0}&\mathbf{0}&\mathbf{0}&\mathbf{0}&\mathbf{0}&\mathbf{0}&\mathbf{0}&\mathbf{H}_{\mathrm{NR}}^{\text{elas}}\\[3.01389pt] \mathbf{0}&\mathbf{0}&\mathbf{0}&\mathbf{0}&\mathbf{0}&\mathbf{0}&\mathbf{0}&\mathbf{0}&\mathbf{0}&\mathbf{H}_{\mathrm{NR}}^{\text{visc}}\\[3.01389pt] \mathbf{H}_{\mathtt{R}_{i}}^{\text{photo}}&\mathbf{0}&\mathbf{H}_{\mathbf{p}_{i}}^{\text{photo}}&\mathbf{H}_{\mathtt{R}_{j}}^{\text{photo}}&\mathbf{0}&\mathbf{H}_{\mathbf{p}_{j}}^{\text{photo}}&\mathbf{0}&\mathbf{0}&\mathbf{0}&\mathbf{H}_{\mathrm{NR}}^{\text{photo}}\\[3.01389pt] \mathbf{0}&\mathbf{0}&\mathbf{0}&\mathbf{0}&\mathbf{0}&\mathbf{0}&\mathbf{I}_{3}&\mathbf{0}&\mathbf{0}&\mathbf{0}\\[3.01389pt] \mathbf{0}&\mathbf{0}&\mathbf{0}&\mathbf{0}&\mathbf{0}&\mathbf{0}&\mathbf{0}&\mathbf{I}_{3}&\mathbf{0}&\mathbf{0}\\[3.01389pt] \mathbf{0}&-\tfrac{1}{\Delta t_{ij}}\mathbf{I}_{3}&\mathbf{0}&\mathbf{0}&\tfrac{1}{\Delta t_{ij}}\mathbf{I}_{3}&\mathbf{0}&\mathbf{0}&\mathbf{0}&-\|\mathbf{g}\|\,\mathbf{J}_{\mathbb{S}^{2}}&\mathbf{0}\par\end{array}\right]$
@@
Elastic term. The elastic prior prevents unrealistic stretching or compression of the graph. If the object bends slightly, nearby nodes may move, but their spacing should not deviate excessively from the reference configuration. This behaviour is encouraged by penalizing changes in pairwise distances:
@@
$\mathcal{L}_{ij,\mathrm{elas}}^{t}=k\,\frac{\left(d_{ij}^{t}-d_{ij}^{0}\right)^{2}}{d_{ij}^{0}},\qquad d_{ij}^{t}=\bigl\lVert\mathbf{x}_{i}^{t}-\mathbf{x}_{j}^{t}\bigr\rVert.$ (12)
In intuitive terms, if nodes preserve their relative spacing the penalty is small; if they stretch or compress significantly, the penalty increases.
@@
Viscous term. While the elastic term controls the shape, we also want to regularize the motion of neighboring nodes over time. Let $\mathbf{s}_{i}^{t}=\mathbf{x}_{i}^{t}-\mathbf{x}_{i}^{t-1}$ be the displacement of node $i$ between two consecutive keyframes. Following [29], we encourage nearby nodes to move in a similar way:
@@
$\mathcal{L}_{ij,\mathrm{visc}}^{t}=b_{ij}\,\bigl\lVert\mathbf{s}_{i}^{t}-\mathbf{s}_{j}^{t}\bigr\rVert^{2},$ (13)
where proximity is encoded by the spatially decaying weights
$b_{ij}=\exp\!\left(-\frac{\left\lVert\mathbf{x}_{i}^{0}-\mathbf{x}_{j}^{0}\right\rVert^{2}}{2\sigma^{2}}\right).$ (14)
Nearby nodes therefore have a strong temporal coupling, promoting smooth and coherent deformations, while distant nodes influence each other only weakly.
@@
Photometric term. Beyond geometric constraints, the formulation also exploits image intensity information. A semi-direct strategy is adopted, in which photometric data association is carried out on Shi–Tomasi features using the modified multi-scale Lucas–Kanade tracker introduced in [30]. When a deformation node is visible in both keyframes t1t{-}1 and $t$, its image projections are expected to correspond to pixels with similar intensities, following the classical brightness constancy assumption:
$\mathcal{L}_{i,\mathrm{photo}}^{t}=\Big(I^{t}(\mathbf{u}_{i}^{t})-\alpha_{i}I^{t-1}(\mathbf{u}_{i}^{t-1})+\beta_{i}\Big)^{2},$ (15)
@@
where $\mathbf{u}_{i}^{t}$ is obtained by projecting $\mathbf{x}_{i}^{t}$ into the image and $I^{t}(\cdot)$ is evaluated via bilinear sampling. Also a local illumination invariance is achieved by computing local gain $\alpha_{i}$ and bias $\beta_{i}$ terms for each point. This term encourages nodes to move consistently with the apparent texture motion in the images.
Combining the previous components yields the full non-rigid regularization for keyframe pair (t1,t)(t{-}1,t):
$\mathcal{L}_{\mathrm{NR}}^{t}=\sum_{(i,j)\in\mathcal{E}}\!\Big(\mathcal{L}_{ij,\mathrm{elas}}^{t}+\mathcal{L}_{ij,\mathrm{visc}}^{t}\Big)+\sum_{i\in\mathcal{V}}\mathcal{L}_{i,\mathrm{photo}}^{t},$ (16)
with $\mathcal{V}$ the set of deformation nodes.
IV-E Overall cost function
@@
The visual–inertial estimation problem is formulated as a nonlinear least-squares optimization over a fixed-size sliding window of keyframes. The window contains the most recent $N$ keyframes and defines the set of state variables over which all measurement constraints are applied. Within this window, the system jointly optimizes rigid-body motion, inertial parameters, and non-rigid deformation variables.
@@
Let $\mathcal{W}=\{t_{0},t_{0}{+}1,\dots,t_{0}{+}N{-}1\}$ denote the active sliding window. As new keyframes are added, the oldest keyframe $t_{0}$ is removed from the window and its corresponding state variables are marginalized. This marginalization step condenses past information into a compact prior term $\mathcal{L}_{\mathrm{prior}}$, which preserves estimator consistency while keeping the computational complexity bounded.
Following [31], only consecutive temporal pairs $\{t{-}1,t\}$ with $t\in\mathcal{W}\setminus\{t_{0}\}$ are considered. For each pair, inertial preintegration, geometric reprojection, and non-rigid regularization terms are accumulated. The resulting objective function minimized by the system is given by
$\mathcal{L}=\sum_{t\in\mathcal{W}\setminus\{t_{0}\}}\Big(\mathcal{L}_{\mathrm{imu}}^{t}+\mathcal{L}_{\mathrm{vision}}^{t}+\lambda_{\mathrm{NR}}\,\mathcal{L}_{\mathrm{NR}}^{t}\Big)\,+\mathcal{L}_{prior},$ (17)
@@
where $\mathcal{L}_{\mathrm{imu}}^{t}$, $\mathcal{L}_{\mathrm{vision}}^{t}$, and $\mathcal{L}_{\mathrm{NR}}^{t}$ are defined in (9), (10), and (16), respectively. The term $\mathcal{L}_{\mathrm{prior}}$ represents the prior induced by marginalization, while the scalar $\lambda_{\mathrm{NR}}$ balances the non-rigid regularization against the visual and inertial contributions.
IV-F Observability analysis
@@
Following [27], observability is assessed from a discrete-time perspective by linearizing the measurement residuals over a finite temporal window. Although Lie derivatives are not explicitly computed, the resulting Jacobians capture the local sensitivity of the measurements to the state in a manner analogous to the continuous-time formulation. As shown in [32, 15], the rank of the stacked Jacobian matrix determines the locally observable directions and is critical for estimator consistency.
@@
Within this framework, observability is analyzed over a two-keyframe window $\{t{-}1,t\}$, jointly considering inertial preintegration, geometric visual, and non-rigid deformation residuals acting on the state (4). The corresponding observability matrix $\mathcal{O}$ is obtained by stacking the Jacobians of all residuals with respect to the state vector $\boldsymbol{\xi}$, yielding
@@
$\mathcal{O}=\begin{aligned} \Big[&\frac{\partial\mathbf{r}_{\Delta\mathtt{R}}^{t}}{\partial\boldsymbol{\xi}}^{\top},\frac{\partial\mathbf{r}_{\Delta\mathbf{v}}^{t}}{\partial\boldsymbol{\xi}}^{\top},\frac{\partial\mathbf{r}_{\Delta\mathbf{p}}^{t}}{\partial\boldsymbol{\xi}}^{\top},\frac{\partial\mathbf{r}_{\mathrm{vision}}^{t}}{\partial\boldsymbol{\xi}}^{\top},\frac{\partial\mathbf{r}_{\mathrm{elas}}^{t}}{\partial\boldsymbol{\xi}}^{\top},\\ &\hskip 16.38895pt\hskip 16.38895pt\frac{\partial\mathbf{r}_{\mathrm{visc}}^{t}}{\partial\boldsymbol{\xi}}^{\top},\frac{\partial\mathbf{r}_{\mathrm{photo}}^{t}}{\partial\boldsymbol{\xi}}^{\top},\frac{\partial\mathbf{r}_{\mathbf{b}^{g}}^{t}}{\partial\boldsymbol{\xi}}^{\top},\frac{\partial\mathbf{r}_{\mathbf{b}^{a}}^{t}}{\partial\boldsymbol{\xi}}^{\top},\frac{\partial\mathbf{r}_{\mathbf{g}}^{t}}{\partial\boldsymbol{\xi}}^{\top}\Big].\end{aligned}$ (18)
@@
TABLE II: Results on the Drunkard’s Dataset. We report mean ATE RMSE, translational RPE, and average number of successfully tracked frames per scene. Each row corresponds to the average performance over all scenes at the same deformation level. Best result per metric in bold. Last row represents the mean for each column.

### v2

@@ III-C Reprojection error
$\mathcal{L}_{\mathrm{rep}}^{\tau}=\sum_{i\in\mathcal{P}}\left\|\mathbf{z}_{i}^{\tau}-\pi\!\left(\mathtt{R}_{\tau},\mathbf{t}_{\tau},\mathbf{x}_{i}^{\tau}\right)\right\|_{\boldsymbol{\Sigma}_{i}^{\tau}},$ (10)
where $\mathbf{z}_{i}^{\tau}\in\Omega$, $\Omega$ being the image domain, denotes the image measurement of the point feature $\mathbf{x}_{i}^{\tau}\in\mathbb{R}^{3}$ at time $t_{\tau}$, $\boldsymbol{\Sigma}_{i}^{\tau}$ is its associated covariance [28] and $\pi(\cdot)$ the camera projection model. We will denote as $\mathcal{P}=\{i\}_{i=1}^{M}$ the full set of 3D points tracked. Note that, differently from rigid visual-inertial odometry, the 3D scene points $\mathbf{x}_{i}^{\tau}$ may change their position at each temporal instant $\tau$.
III-D Non-rigid residual
@@
We will assume that our imaged scenes may exhibit non-rigid motions, such as those exhibited by deformable clothes or cables under physical interaction. We will model these effects by a lightweight deformation graph that we build over a subset $\mathcal{D}=\{i\}_{i=1}^{D}\subset\mathcal{P}$ composed by the $D$ longer feature tracks from $\mathcal{P}$. The tracked points $\mathbf{x}_{i}^{\tau}$ define the nodes of the graph. We will define the reference keyframe as the first one at $\tau=1$. Two nodes $i,j\in\mathcal{D}$ are connected by an edge $(i,j)$ if their distance in the reference keyframe is below a spatial threshold. This provides a compact neighborhood structure that encodes how nearby points of the scene relate to each other. The reference distance between nodes is defined in the reference keyframe as
$d_{ij}^{1}=\bigl\lVert\mathbf{x}_{i}^{1}-\mathbf{x}_{j}^{1}\bigr\rVert.$ (11)
The full non–rigid residual is built up by several terms that we detail next.
@@
$\mathcal{O}=\left[\begin{array}[]{cccccccccc}-\mathbf{J}_{r}^{-1}(\boldsymbol{\phi})\mathtt{R}_{\tau+1}^{\top}\mathtt{R}_{\tau}&\mathbf{0}&\mathbf{0}&\mathbf{J}_{r}^{-1}(\boldsymbol{\phi})&\mathbf{0}&\mathbf{0}&\mathbf{0}&\mathbf{0}&\mathbf{0}&\mathbf{0}\\[3.01389pt] \left(\mathtt{R}_{\tau}^{\top}(\mathbf{v}_{\tau+1}-\mathbf{v}_{\tau}-\mathbf{g}\Delta T)\right)^{\wedge}&-\mathtt{R}_{\tau}^{\top}&\mathbf{0}&\mathbf{0}&\mathtt{R}_{\tau}^{\top}&\mathbf{0}&\mathbf{0}&-\mathtt{R}_{\tau}^{\top}\Delta T&-\|\mathbf{g}\|\mathtt{R}_{\tau}^{\top}\Delta T&\mathbf{0}\\[3.01389pt] \left(\mathtt{R}_{\tau}^{\top}(\mathbf{t}_{\tau+1}-\mathbf{t}_{\tau}-\mathbf{v}_{\tau}\Delta T-\tfrac{1}{2}\mathbf{g}\Delta T^{2})\right)^{\wedge}&-\mathtt{R}_{\tau}^{\top}\Delta T&-\mathtt{R}_{\tau}^{\top}&\mathbf{0}&\mathbf{0}&\mathtt{R}_{\tau}^{\top}&\mathbf{0}&-\mathtt{R}_{\tau}^{\top}\tfrac{1}{2}\Delta T^{2}&-\tfrac{1}{2}\|\mathbf{g}\|\mathtt{R}_{\tau}^{\top}\Delta T^{2}&\mathbf{0}\\[3.01389pt] \mathbf{H}_{\mathtt{R}_{\tau}}^{\text{vis}}&\mathbf{0}&\mathbf{H}_{\mathbf{t}_{\tau}}^{\text{vis}}&\mathbf{H}_{\mathtt{R}_{\tau+1}}^{\text{vis}}&\mathbf{0}&\mathbf{H}_{\mathbf{t}_{\tau+1}}^{\text{vis}}&\mathbf{0}&\mathbf{0}&\mathbf{0}&\mathbf{H}_{\mathrm{nr}}^{\text{vis}}\\[3.01389pt] \mathbf{0}&\mathbf{0}&\mathbf{0}&\mathbf{0}&\mathbf{0}&\mathbf{0}&\mathbf{0}&\mathbf{0}&\mathbf{0}&\mathbf{H}_{\mathrm{nr}}^{\text{elas}}\\[3.01389pt] \mathbf{0}&\mathbf{0}&\mathbf{0}&\mathbf{0}&\mathbf{0}&\mathbf{0}&\mathbf{0}&\mathbf{0}&\mathbf{0}&\mathbf{H}_{\mathrm{nr}}^{\text{visc}}\\[3.01389pt] \mathbf{H}_{\mathtt{R}_{\tau}}^{\text{photo}}&\mathbf{0}&\mathbf{H}_{\mathbf{t}_{\tau}}^{\text{photo}}&\mathbf{H}_{\mathtt{R}_{\tau+1}}^{\text{photo}}&\mathbf{0}&\mathbf{H}_{\mathbf{t}_{\tau+1}}^{\text{photo}}&\mathbf{0}&\mathbf{0}&\mathbf{0}&\mathbf{H}_{\mathrm{nr}}^{\text{photo}}\\[3.01389pt] \mathbf{0}&\mathbf{0}&\mathbf{0}&\mathbf{0}&\mathbf{0}&\mathbf{0}&\mathbf{I}_{3}&\mathbf{0}&\mathbf{0}&\mathbf{0}\\[3.01389pt] \mathbf{0}&\mathbf{0}&\mathbf{0}&\mathbf{0}&\mathbf{0}&\mathbf{0}&\mathbf{0}&\mathbf{I}_{3}&\mathbf{0}&\mathbf{0}\\[3.01389pt] \mathbf{0}&-\tfrac{1}{\Delta T}\mathbf{I}_{3}&\mathbf{0}&\mathbf{0}&\tfrac{1}{\Delta T}\mathbf{I}_{3}&\mathbf{0}&\mathbf{0}&\mathbf{0}&-\|\mathbf{g}\|\,\mathbf{J}_{\mathbb{S}^{2}}&\mathbf{0}\par\end{array}\right]$
@@
Elastic constraint. We will assume an elastic relation between points in the deformation graph, preventing unrealistic stretching or compression. If part of the deformable object bends, nearby nodes may also experience non-rigid deformation, but their spacing should not deviate excessively from the reference configuration. We promote this behavior by penalizing changes in pairwise distances.
@@
$\mathcal{L}_{ij,\mathrm{elas}}^{\tau}=\kappa\,\frac{\left(d_{ij}^{\tau}-d_{ij}^{1}\right)^{2}}{d_{ij}^{1}},\qquad d_{ij}^{\tau}=\bigl\lVert\mathbf{x}_{i}^{\tau}-\mathbf{x}_{j}^{\tau}\bigr\rVert.$ (12)
where $\kappa$ is an elastic deformation constant. In words, this term will penalize excessive stretching or compression, so small deformations will be preferred for the same IMU and multi-view constraints.
@@
Viscous constraint. While the elastic term controls shape deviations between times $t_{\tau}$ and $t_{1}$, we additionally impose constraints on the kinematics via a viscous term. Let $\mathbf{s}_{i}^{\tau}=\mathbf{x}_{i}^{\tau}-\mathbf{x}_{i}^{\tau-1}$ be the displacement of node $i$ between two consecutive keyframes. Following [29], we encourage nearby nodes to move in a similar manner
@@
$\mathcal{L}_{ij,\mathrm{visc}}^{\tau}=b_{ij}\,\bigl\lVert\mathbf{s}_{i}^{\tau}-\mathbf{s}_{j}^{\tau}\bigr\rVert^{2},$ (13)
where proximity is encoded by spatially decaying weights $b_{ij}$, which impose strong temporal coupling between spatially close frames, promoting smooth and coherent deformations
$b_{ij}=\exp\!\left(-\frac{\left\lVert\mathbf{x}_{i}^{1}-\mathbf{x}_{j}^{1}\right\rVert^{2}}{2\sigma^{2}}\right),$ (14)
being $\sigma$ a constant modeling the strength of the viscous constraint.
@@
Photometric constraint. In addition to geometric reprojection error, we also leverage image intensity information. A semi-direct strategy is adopted, in which photometric data association is carried out on Shi–Tomasi features using the modified multi-scale Lucas–Kanade tracker introduced in [30]. For every deformation node, its image projections in consecutive keyframes $\tau{-}1$ and $\tau$ are expected to correspond to pixels with similar intensities, following the classical brightness constancy assumption:
$\mathcal{L}_{i,\mathrm{photo}}^{\tau}=\Big(I^{\tau}(\mathbf{u}_{i}^{\tau})-\alpha_{i}I^{\tau-1}(\mathbf{u}_{i}^{\tau-1})+\beta_{i}\Big)^{2},$ (15)
@@
where $\mathbf{u}_{i}^{\tau}=\pi\left(\mathtt{R}_{\tau},\mathbf{t}_{\tau},\mathbf{x}_{i}^{\tau}\right)$ and $I^{\tau}(\mathbf{u}_{i})$ is evaluated via bilinear sampling. Also a local illumination invariance is achieved by computing local gain $\alpha_{i}$ and bias $\beta_{i}$ terms for each tracked image point. This term encourages nodes to move consistently with the apparent texture motion in the images.
The combination of the elastic, viscous and photometric terms yields the full non-rigid residual for a keyframe pair $(\tau{-}1,\tau)$:
$\mathcal{L}_{\mathrm{nr}}^{\tau}=\sum_{(i,j)\in\mathcal{E}}\!\Big(\mathcal{L}_{ij,\mathrm{elas}}^{\tau}+\mathcal{L}_{ij,\mathrm{visc}}^{\tau}\Big)+\sum_{i\in\mathcal{D}}\mathcal{L}_{i,\mathrm{photo}}^{\tau}$ (16)
where $\mathcal{E}$ is the set of all edges in the deformation graph.
III-E Overall cost function
@@
The visual–inertial estimation problem is formulated as a nonlinear least-squares optimization over a fixed-size sliding window $\mathcal{W}$. The window contains the most recent $N$ keyframes and defines the state vector $\xi$ over which constraints are applied. Within $\mathcal{W}$, our optimization jointly optimizes rigid-body motion, inertial parameters and non-rigid deformation variables as defined in Eq. 1.
@@
As the window slides, new keyframes are added and the oldest ones are removed by marginalizing their corresponding states. This marginalization step condenses past information into a compact prior term $\mathcal{L}_{\mathrm{prior}}$, which preserves some degree of consistency while keeping the computational complexity bounded. The resulting objective function minimized by the system is given by
$\mathcal{L}=\sum_{\tau\in\mathcal{W}}\Big(\mathcal{L}_{\mathrm{imu}}^{\tau}+\mathcal{L}_{\mathrm{rep}}^{\tau}+\lambda_{\mathrm{nr}}\,\mathcal{L}_{\mathrm{nr}}^{\tau}\Big)\,+\mathcal{L}_{\mathrm{prior}},$ (17)
@@
where the scalar $\lambda_{\mathrm{nr}}$ balances the non-rigid regularization against visual and inertial terms.
III-F Observability analysis
@@
Following [27], observability is assessed from a discrete-time perspective by linearizing the measurement residuals over a finite temporal window. Although Lie derivatives are not explicitly computed, the resulting Jacobians capture the local sensitivity of the residuals with respect to the state, analogously to the continuous-time formulation. As shown in [31, 13], the rank of the stacked Jacobian matrix determines the locally observable directions and is critical for estimator consistency.
@@
Within this framework, we analyze observability over a two-keyframe window $\{\tau{-}1,\tau\}$, jointly considering inertial preintegration, geometric visual, and non-rigid deformation residuals acting on the state (1). The corresponding observability matrix $\mathcal{O}$ is obtained by stacking the Jacobians of all residuals with respect to the state vector $\boldsymbol{\xi}$, yielding
@@
$\mathcal{O}=\Big[$ $\frac{\partial\mathbf{r}_{\Delta\mathtt{R}}^{\tau}}{\partial\boldsymbol{\xi}}^{\top},\frac{\partial\mathbf{r}_{\Delta\mathbf{v}}^{\tau}}{\partial\boldsymbol{\xi}}^{\top},\frac{\partial\mathbf{r}_{\Delta\mathbf{t}}^{\tau}}{\partial\boldsymbol{\xi}}^{\top},\frac{\partial\mathbf{r}_{\mathrm{rep}}^{\tau}}{\partial\boldsymbol{\xi}}^{\top},\frac{\partial\mathbf{r}_{\mathrm{elas}}^{\tau}}{\partial\boldsymbol{\xi}}^{\top},$ (18)
$\hskip 18.49988pt\frac{\partial\mathbf{r}_{\mathrm{visc}}^{\tau}}{\partial\boldsymbol{\xi}}^{\top},\frac{\partial\mathbf{r}_{\mathrm{photo}}^{\tau}}{\partial\boldsymbol{\xi}}^{\top},\frac{\partial\mathbf{r}_{\mathbf{b}^{g}}^{\tau}}{\partial\boldsymbol{\xi}}^{\top},\frac{\partial\mathbf{r}_{\mathbf{b}^{a}}^{\tau}}{\partial\boldsymbol{\xi}}^{\top},\frac{\partial\mathbf{r}_{\mathbf{g}}^{\tau}}{\partial\boldsymbol{\xi}}^{\top}\Big].$
@@
TABLE I: Drunkard’s Dataset results. We report ATE RMSE, RPE, and number of successfully tracked frames. Each row averages all scenes at the same deformation level. Best result per metric in bold. Last row shows per-column means.